An iron ball of mass 3\,\,kg3kg is hanging at rest tied to the roof by a vertical massless string of length 10m. Then a horizontal force F=14\,\,NF=14N acts on the ball till the ball reaches a point where the string makes an angle of 37 with the vertical. Find the speed of the ball at this instant. (Assume g=10m/s2 sin37=0.6 )
1
Expert's answer
2020-01-29T14:08:18-0500


In this case,"h=l-l\ cos\theta=l-l\ cos 37\degree=\frac{l}{5}"

As the gravity is acting opposite to direction of ball,

Net work due to gravity"=-(mgh)=-mg\frac{l}{5}"

Net work done due to applied force"=F\times s=F\times l\ sin\theta"

As per work energy theorem,

Change in Kinetic energy=Net work done

"\frac{1}{2}mv^2=Fl\ sin\theta-mg\frac{l}{5}"

Substituting the values,"\theta=37\degree,F=14N,l=10 m,m=3\ Kg" ,

"v=4m\/sec"